OpenStudy (anonymous):

MEDALS AWARDED!!!!! Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -8, and 2 + 3i

OpenStudy (debbieg):

Each root c corresponds to a factor, (x-c) Also, for any complex root a + bi, the complex conjugate a - bi must also be a root. So that means the 3 roots you're given REALLY tell you FOUR roots.... so line up those 4 corresponding factors: (x-4)(x+8)[x-(2 + 3i)][x-(2 - 3i)] And multiply it out.

OpenStudy (anonymous):

yeah, you can multiply out, but there are two easier ways to find the quadratic with zeros \(2+3i\) and \(2-3i\)

OpenStudy (anonymous):

one way is to work backwards set \[x=2+3i\] subtract \(2\) and get \(x-2=3i\) square to get \[(x-2)^2=-9\] or ' \[x^2-4x+4=-9\] then add 9 to get \[x^2-4x+13\] as your quadratic

OpenStudy (anonymous):

the real real easy way, although it requires memorization, is to know that if the quadratic has zeros \(a+bi\) and \(a-bi\) then it is \[x^2-2ab+(a^2+b^2)\] but you have to memorize that it is not hard though and would save you lots of time on a quiz or test

OpenStudy (anonymous):

oops that was wrong it is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (anonymous):

so in your example since \(a+bi=2+3i\) you get \[x^2-2\times 2x+(2^2+3^2)=x^2-4x+13\]

OpenStudy (debbieg):

Nice, that's slick! Great tip. I know that you can make the complex factors easier to multiply by rewriting as a product of sum and difference: [x-(2 + 3i)][x-(2 - 3i)]= [(x-2) - 3i)][(x-2) + 3i)] Then you get: \((x-2)^2 - (3i)^2=x^2-4x+4-(-9)=x^2-4+13\) which is kind of a similar idea, but I think I might like the "backwards" way that @satellite73 showed. :)

OpenStudy (debbieg):

.... even better, I mean. lol

OpenStudy (anonymous):

pretty much the exact same computation right?

OpenStudy (anonymous):

i.e. you end up squaring \(x-2\) and then adding \(9\)

OpenStudy (anonymous):

thank you i got it!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!